Circular No. 5772

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET REINMUTH 2 (1993g) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the 0.9-m Spacewatch telescope at Kitt Peak:

1993 UT  α  δ_{2000}  m_{1}
Feb. 26.45957  15^{h}07^{m}59.^{s}67  −25^{o}33′44′′.7  21.9
26.48942  15 08 00.14  −25 33 50.1  22.0
26.51786  15 08 00.26  −25 33 56.0  21.7
Apr. 21.39013  14 52 12.02  −26 03 41.1
21.39985  14 52 11.63  −26 03 39.5  21.7
21.41247  14 52 10.94  −26 03 37.7  21.7
22.36974  14 51 26.83  −26 01 21.4  20.8
22.37880  14 51 26.49  −26 01 20.8  21.5
22.39174  14 51 25.73  −26 01 18.7  21.3
22.40375  14 51 25.24  −26 01 16.2  22.1

The images were right at the telescope limit in good seeing on Feb. 26. The comet was rather low on Apr. 21, and there was no convincing diffuseness beyond the poor seeing, although the images seemed softer than those of objects of similar brightness nearby. The position is almost precisely in accordance with the prediction on MPC 18258 (ephemeris on MPC 21812).

SUPERNOVAE 1989ab AND 1989ac J. Mueller reports her discovery of two apparent supernovae on plates

taken during the course of the second Palomar Sky Survey with the 1.2-m Oschin Schmidt Telescope. SN 1989ab was found in an anonymous galaxy whose nucleus is located at α = 18^{h}36^{m}45^{s}.06, δ = +32^{o}54′55′′.5 (equinox 1950.0); the supernova is located 13′′ west and 11′′ south of the nucleus and appears at mag ∼ 18 on the discovery plate, which was taken 1989 July 6 by C. Brewer and Mueller. SN 1989ab was also found on a survey plate taken by Brewer, Mueller, and J. D. Mendenhall on 1989 June 30, and it does not appear on original Palomar Sky Survey prints or on a survey J plate taken by Brewer and Mendenhall on 1991 July 15. SN 1989ac is located 20′′ west and 6′′.7 south of the nucleus of UGC 1867 (α = 2^{h}22^{m}.0, δ = +45^{o}15′, equinox 1950.0); the supernova appeared at blue mag ∼ 17.5 on the discovery plate, which was taken 1989 Aug. 31 by Mueller. SN 1989ac was confirmed on a survey plate taken 1989 Sept. 4 by Mendenhall and Mueller, and it does not appear on original Sky Survey prints.

1993 April 22 Daniel W. E. Green